OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (michele_laino):

please note that your sequence is a geometric progression whose first term is -8 and constant = 5. So your first six terms, are: \[a _{1}=-8,a _{2}=-8*5=-40,a _{3}=-5*(-40)=-200\]

OpenStudy (michele_laino):

\[a _{4}=5*(-200)=-1000,a _{5}=5*(-1000)=-5000\] \[a _{6}=5*(-5000)=-25000\]

OpenStudy (anonymous):

oh wow, I get it now ... thank you !!!!!

OpenStudy (anonymous):

can you help me on this one also , because it is different

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. 8, 6, 4, 2, ...

OpenStudy (michele_laino):

in general, we have, for a generic n: \[a _{n}=a _{1}*5^{(n-1)}\]

OpenStudy (anonymous):

I have to make the equation myself now

OpenStudy (anonymous):

would it be 8 + -2(n)

OpenStudy (michele_laino):

now, we have an arthmetic sequence, whose first term is 2, and constant=2, so your f n-term is: \[a _{n}=a _{1}+(n-1)*c=2+(n-1)*2=2+2n-2=2n\]

OpenStudy (anonymous):

sorry , ok I figured that one out myself , I have this question Find the first six terms of the sequence. a1 = -4, an = an-1 + 7 can you show me in steps how I can solve it I am a little confused

OpenStudy (michele_laino):

in this case you have an arithmetic sequence, whose first term is -4, and constant=7. from your recurrence formula, I can write: \[a _{2}=a _{1}+7=-4+7=3\] \[a _{3}=a _{2}+7=3+7=10\]

OpenStudy (michele_laino):

\[a _{4}=a _{3}+7=10+7=17\] finally: \[a _{5}=a _{4}+7=17+7=24\]

OpenStudy (anonymous):

but the answer choices are an = -3 - 2 an = -3 + -2(n) an = -3 + -2(n + 1) an = -3 + -2(n - 1)

OpenStudy (michele_laino):

for generic nth term we have, applying the known formula about aruthmetic sequences: \[a _{n}=a _{1}+(n-1)*c=-4+(n-1)*7=\]

OpenStudy (anonymous):

is it c ?

OpenStudy (michele_laino):

\[=-4+7n-7=7n-11=a_{n}\]

OpenStudy (anonymous):

that is not one of the choices :(

OpenStudy (michele_laino):

I see!

OpenStudy (michele_laino):

even if my formula is correct!

OpenStudy (anonymous):

so what do I put

OpenStudy (michele_laino):

sorry, what is the meaning of: an = -3 + -2(n - 1)

OpenStudy (michele_laino):

maybe a number is missing between +sign and -sign?

OpenStudy (anonymous):

it really means -3 + (-2)(n-1)

OpenStudy (michele_laino):

I'm very sorry, I don't know. Nevertheless I can say you that if your formulas, namely: a1 = -4, an = an-1 + 7 are correct, then the expression for the n-th term is the one that I wrote above, namely: a_n=7n-11